# DiVerify: Hardening Identity-Based Software Signing with Programmable Diverse-Context Scopes

### v2

@@ 3.2 Threat Model
We describe three main types of adversaries with the following capabilities:
  1. 1.
@@
  3. 3.
    Type III: Compromised Signing Tool. The attacker can tamper with the client software to sign software beyond what the user originally intended, allowing unauthorized code to be signed.
Examples of each type
Recent events illustrate:
  1. 1.
    ESLint (Type I: Credential Compromise): ESLint, a widely used JavaScript linter, was compromised in 2018 when attackers used stolen npm credentials to upload malicious package [92]. Because the system relied only on token-based authentication, it could not tell legitimate actions from malicious ones.
  2. 2.
    Midnight Blizzard (Type II: IdP Compromise): In 2023, the nation-state actor Midnight Blizzard compromised a legacy Microsoft test account to create malicious OAuth apps with elevated privileges [55]. These apps granted persistent access to Exchange Online, bypassing identity checks. Since the tokens were issued by a trusted provider and used valid credentials, the system lacked the context to detect the compromise.
  3. 3.
    Codecov (Type III: Tool Compromise): A widely used code coverage tool integrated into CI was compromised in 2021, when attackers added code to exfiltrate sensitive information [10]. The attack persisted undetected as users implicitly trusted the tool, with no way to verify its runtime state. Other previous attacks have also taken this form of abuse [85, 40, 44].
Assumptions
We assume legitimate signers do not sign untrusted code, verifiers properly check all signatures, the TEE setup remains secure, and not all identity providers can be compromised at the same time. Rationales and additional assumptions are discussed in Appendix˜A.
4 System Goals and Overview
This section shows our security and deployment goals and gives an overview of DiVerify’s approach.
4.1 System Goals
In order to address the attackers described in our Threat Model, DiVerify aims to extend authentication beyond traditional credentials; harden the security posture of signing clients; and limit reliance on any single identity provider by requiring multiple, independent sources of verification. We formalize these notions into goals:
Security Goals:
@@ 4.1 System Goals
    Resilience Against Identity Theft: Signing events are allowed only after verification from multiple independent sources, which limits trust assumptions on identity providers. This way, if an attacker compromises a Signer’s credentials, an attacker is unable to complete a signer request without a collection of these.
@@
    Resilience Against IdP compromise: In addition to diversity of credentials, a Signer can request multiple identity providers to prove a diverse identity claim. This ensures that a compromised provider cannot impersonate its user unless it has also compromised a threshold number of other independent providers.
@@
Deployment Goals:
In addition to the security goals, DiVerify is also designed to facilitate adoption within existing software supply chain tools:
  1. D1
    Backwards compatibility in a mixed model within the code signing ecosystems
  2. D2
    Little-to-no changes to existing tooling
4.2 System Overview
DiVerify follows a disaggregated trust model where trust in a signature is subject to a combination of independently verifiable claims about identity, context, and delegated trust. In practice, this manifests in a similar way as the identity-based software signing flow described above. However, in contrast, we rely on a generalized party, a scope provider, which replaces the role of an IdP. In order to dis-aggregate trust, a set of Scope Providers will provide individual proofs regarding the Signer’s properties.
Figure 2 presents a high-level view of the system. In the same way as a regular identity-based software signing setting, the flow initiates when a Signer requests that a client performs the signing process. Instead of requesting an identity token from an IdP, it will query a set of scope providers to obtain multiple scopes. These scopes can include, e.g., OIDC identity token as proof of identity, a device fingerprint or OS attestation as proof of environment, and a hardware security token proof as a second factor. In parallel, the client also collects integrity evidence about the signing environment itself.
Having collected these scopes, the client must aggregate them into a DiVerify proof which shows how the signer was authenticated, and also embed a proof of its own integrity. This can be carried out in various ways such as using TPM-ME and a seal/unseal system [30], or by moving a portion of the client code into a separate party inside a TEE. The resulting DiVerify proof serves a role similar to an identity-based signing certificate, but with richer information that a verifier can use to assess whether they trust a particular signature over a piece of software. Lastly, the verifier is responsible for validating these proofs and verifying the resulting signatures.
Figure 2: DiVerify Overview. DiVerify extends traditional identity providers into scope providers, which assert claims about the signer. The DiVerify daemon, running in a secure enclave, coordinates and proves its own integrity to a verifier.
For example, consider a flow where a signer Alice wants to sign an artifact in a DiVerify setting. They want verifiers to know that not only are they alice@software.sh, but also that they signed their software inside her MacBook Pro (identified with the fingerprint 0xCAFEC0FFEE), and using a second factor from her Yubikey. Users of Alice’s software are aware of this, and will use a DiVerify policy to agree that valid software from Alice needs to contain proofs of the above.
Alice will then follow this flow to generate the signature, which Bob can then validate:
  1. 1.
    Alice will issue a signing request through her client tool, specifying a trust level (i.e., the set of properties described above).
  2. 2.
    The client acting on behalf of the Alice prepares a signing request object, determines which scopes meet the trust level and prompts the her to authenticate with the corresponding scope providers.
  3. 3.
    For each scope requested, each provider will issue signed scope to the client if Alice is successful. In this particular case, Alice authenticates against an identity provider to prove she is indeed alice@software.sh, she obtains a fingerprint from her operating system non-interactively, and touches her Yubikey to generate a second factor scope signature.
  4. 4.
    The client verifies the returned scopes, signs the software, and appends a last integrity check to prove its own integrity (e.g., by creating a TEE remote-attestation object). Then it aggregates these scopes and the attestation into a DiVerify proof.
  5. 5.
    The signed artifact, along with its DiVerify proof, is submitted to a package repository for storage and distribution.
  6. 6.
    Bob retrieves the artifact, extracts its associated policy, and checks that: (a) the signature is valid and it applies to the obtained artifact. (b) the provided proof satisfies the agreed upon policy (that describes which scopes are needed and their expected values). (c) the proof was generated by a legitimate and trustworthy client software.
[…]
Of note, some of the techniques of DiVerify overlap. For example, the use of contextual scopes will also reduce the likelihood that an attacker is able to sign on behalf of the client using stolen credentials. This is because an attacker must also be able to initiate a signing request in the same context (e.g., hardware) as the Signer.
5 DiVerify Architecture
Having provided an overview of DiVerify , we move on to describe the components of DiVerify , followed by the protocol of how they interact to support secure and verifiable signing.
5.1 Components
@@
manage and issue claims about a signer’s identity, attributes, or environment context. Each operates within a distinct trust domain, that is, an independently managed authority responsible for issuing and validating its own set of claims about signers. They follow a specific protocol to verify properties such as control of an email address, possession of a cryptographic key, or ownership of a device and then issues a signed statement attesting to that property. DiVerify does not propose new protocols for scope providers, but builds on existing authentication standards whenever possible. Scope providers must be able to authenticate themselves to the tool and prevent scope replay to ensure that claims are both trustworthy and fresh. Since not all providers support both of these properties natively, DiVerify introduces strategies to “wrap” them and implement arbitrary scope providers at a meta-protocol level. We discuss these strategies in section˜5.3.
@@
is a signer-facing component responsible for initiating signing requests. This can be instantiated as a command-line tool or library integrated into developer environments. The client operates within the signer’s local context and has access to the artifact to sign. It serves as the point of integration between the signer and the rest of the DiVerify architecture. Because the client runs in a potentially untrusted setting, we split the functionality of a traditional signing client into two parts: a lightweight front-end Client and a privileged Daemon.
@@
is an end-user side component responsible for validating signed artifacts and assessing their trustworthiness. It operates independently of the signing process and has no privileged relationship with the signer. A DiVerify verifier will receive a triple $(K_{p},SW_{a},S_{sw})$ as well as a policy $P_{sw}$ and return either $OK$ if the software is trusted or $ERROR$ if any of the validation steps failed.
5.2 DiVerify Protocol
5.2.1 Signing Request Initialization
DiVerify-enabled signing tools can advertise which scope providers they support, allowing signers to know what factors can be used. We define a trust levels configuration that groups scopes into levels based on the types of claims. Higher trust levels correspond to requiring more diverse claims (an example is shown in Listing LABEL:lst:diverify-config-structure in the Appendix), increasing overall confidence in the signer’s legitimacy. When the DiVerify daemon starts, it establishes channels to interact with the client and with scope providers.
To initialize, a Signer interacts with a Client to initialize a signing request. This can be done by issuing a command through the DiVerify client. The parameters to initialize a signing request include trust level, and an identifier (e.g., a filename, or docker image tag) of the software artifact to sign.
5.2.2 Scope Retrieval and Signing
With a signing request initialized, the client submits a request to the daemon. Following Algorithm 1, the daemon first determines which scopes are required for the requested trust level (line 2). These required scopes are independent, and can be obtained in parallel from their respective providers. The daemon generates a fresh nonce to use in all scope requests to bind them to this signing session. It then iterates over each target scope provider and initiates an authentication protocol to obtain the scope (lines 4–7).
@@ 5.2.2 Scope Retrieval and Signing
1:In: Artifact AA, TrustLevel $L$
2:Requires: Scope Provider $P$, TrustConfig
@@
4:procedure SignRequest(AA, $L$)
5:  $req\_scopes\leftarrow\text{TrustConfig}(L)$
6:  $scopes\leftarrow\{\}$
7:  for $m\in req\_scopes$ do
8:   $r\leftarrow\text{AuthenticateSigner}({P}_{m})$
9:   $scopes[m]\leftarrow r$
10:  end for
11:  if not ValidScope(scopesscopesthen
12:   return Err
@@
14:  $sk,pk=\text{GenerateKeyPair}()$
15:  $S\leftarrow\text{Sign}_{sk}(A)$
16:  $custom\_data\leftarrow\text{Sign}_{sk}(hash(scopes))$
17:  $quote\leftarrow\text{GetQuote}(custom\_data)$
18:  $dvp\leftarrow\text{MakeDiVerifyProof}(scopes,quote,pk)$
19:  return $S$, dvpdvp
20:end procedure
Each request follow a DiVerify specific meta-protocol. The protocol proceeds as follows: first, the daemon submits a request such as $request(n,claim)$. Each provider responds with a tuple $(\mathit{sign(proof)},\ \mathit{proof})$, where $\mathit{proof}$ is a protocol-specific claim (e.g., an oidc token), and the accompanying signature binds it to the daemon’s nonce. Of note, the nonce itself is embedded inside of the proofproof. This could be trivially surfaced as part of the top-level response (and turn the tuple above into a triple), but this would not retain backward compatibility to other natively supported protocols such as OIDC.
On the scope provider side, scopes may request the signer to prove the claimclaim. For example, if a Signer requests an OIDC provider to sign a scope for alice@software.sh, the provider may request a $username,password$ tuple to be input through the client or a browser. Other scope providers may not require user interaction, and will reply with the proof directly.
Once the daemon receives scope responses from providers, it verifies each tuple $(\mathit{sign(proof)},\ \mathit{proof})$ by checking that the signature is valid and that the proof includes the correct daemon-generated nonce (line 8). Since the daemon is trusted, the Signature from the scope provider does not need to be forwarded to the verifier. After successful verification, the daemon proceeds to generate ephemeral signing keys and perform the signing operation.
5.2.3 Attestation Binding
The daemon must prove to a verifier that it is running untampered inside an enclave via remote attestation. To prevent a client from replaying attestations using scopes or signing key it controls, the daemon must cryptographically bind the scope and signing key to the attestation.
To achieve this, the daemon first signs a hash of the validated scopes using the signing key (line 13). It then embeds the resulting signature into the attestation quote (line 14).
Because this field is included in the signed remote attestation quote, it cryptographically binds the signing key to the TEE identity and state.
5.2.4 DiVerify Proof Construction
Since a verifier needs to have a direct view on the scope claims, and key used for signing to allow it to verify a signed artifact, DiVerify needs to be able to provide that information. This is done by providing a DiVerify proof (as shown in Listing LABEL:lst:proof-structure in the Appendix), which consists of the following: the validated scopes (i.e., the data that the daemon validated and signed), the signing public key pair, and the remote attestation. The daemon then bundles the proof and the signature, and send to the client who uploads it to the package repository.
5.2.5 Verification
To verify a signature, a verifier needs the signed artifact, the generated signature, DiVerify proof, and policy. The DiVerify policy defines expected signer claims and a logical rule of how these claims must combine to satisfy trust. We discuss its semantics below and how it is securely distributed using TUF[72] in section˜5.4.
@@ 5.2.5 Verification
The verification process follows the flow outlined in Algorithm 2. The procedure begins with the verifier retrieving the current verification policy from TUF. It then extracts the attestation quote, the associated custom data, the public key, and the scopes from the DiVerify proof. Next, it validates the integrity of the quote (line 7), confirms that the scope digest matches the custom data using the public key (line 10). Then it evaluates whether the scopes comply with the retrieved policy. Finally, it validates that the artifact using the public key. If any of these checks fail, the artifact is deemed untrusted and is consequently rejected.
Algorithm 2 Artifact Verification Flow
1:Input: Artifact A, ArtifactSignature $S$, DiVerify proof dvpdvp
2:Requires: RSTUF $T$
3:Output: Boolean verification result
4:procedure VerifyArtifact(A, S)
5:  policypolicy $\leftarrow$ $T.\text{GetPolicy}()$
6:  quotequote $\leftarrow$ $dvp.\text{quote}$
7:  cutom_datacutom\_data $\leftarrow$ $quote.\text{cutom\_data}$
8:  pkpk $\leftarrow$ $dvp.\text{pubkey}$
9:  scope_digestscope\_digest $\leftarrow$ $hash(dvp.scopes)$
10:  if not ValidQuote(quotequotethen
11:   return false
12:  end if
13:  if not VerifySignature(scope_digestscope\_digest,custom_datacustom\_data,pkpkthen
14:   return false
15:  end if
16:  if not EvaluatePolicy($dvp.scopes$, policypolicythen
17:   return false
18:  end if
19:  if not VerifySignature(AA,$S$,pkpkthen
20:   return false
21:  end if
22:  return true
23:end procedure
Diverify Policy Semantics
We recognize that a legitimate signer may not always be able to satisfy all scopes at every signing instance. For example, a signer using a personal computer may authenticate via an identity provider, prove possession of a hardware token, and demonstrate that signing occurred on a specific, trusted machine fingerprint. However, when operating from a remote environment, the same signer may still authenticate using an identity provider or hardware key—but may no longer be able to prove that signing occurred on the expected, fingerprinted device.
DiVerify supports this flexibility, and allows software maintainers to define trust policies that specify which scopes are required and what constraints must hold for a signature over the software to be accepted.
In the sample policy in listing LABEL:lst:diverify-scope, each field defines an expected value that must match a corresponding claim in the DiVerify proof. The rule (lines 10) field expresses a Boolean condition over these claims, defining the minimal combination required for trust. In the sample, the policy requires the signer to authenticate with the specified identity and provider, and to present either the expected device fingerprint or a hardware token. Additionally, the proof must provide a remote attestation that matches the known trusted signer measurement.
1"identity": {
2 "email": "alice@software.sh",
3 "provider": "https://accounts.google.com",
4"device_fingerprint": {
5 "expected_fpr": "0x1234abc...",
6"signer_measurement": "sha256:5678def...",
7"security_key": {
8 "slot9a_public_key": "Fkw...EwYH",
9 "slotf9_attestation_cert": "MIE..AB",}
10"rule": "(identityANDprovider)AND(device_fingerprintORhardware_key)AND(signer_measurement)"
Listing 1: Sample DiVerify Policy Structure
5.3 Integrating Scope Providers

### v3

@@ 3.2. Threat Model
We describe three main types of adversaries with the following capabilities, reflecting the different ways an adversary can obtain signing authority:
  1. (1)
@@
  3. (3)
    Type III: Compromised Signing Tool. The attacker can tamper with the client software to sign software beyond what the user originally intended, allowing unauthorized signatures.
These attacks are not theoretical, as shown by the events below:
  1. (1)
    ESLint (Type I: Credential Compromise): ESLint, a widely used JavaScript linter, was compromised in 2018 when attackers used stolen npm credentials to upload malicious package (Zhu, 2018). Because the system relied only on token-based authentication, it could not tell legitimate actions from malicious ones.
  2. (2)
    Midnight Blizzard (Type II: IdP Compromise): In 2023, the nation-state actor Midnight Blizzard compromised a legacy Microsoft test account to create malicious OAuth apps with elevated privileges (Microsoft, 2024). These apps granted persistent access to Exchange Online, bypassing identity checks. Since the tokens were issued by a trusted provider, the system lacked the context to detect the compromise.
  3. (3)
    Codecov (Type III: Tool Compromise): A widely used code coverage tool integrated into CI was compromised in 2021, when attackers added code to exfiltrate sensitive information (Codecov, 2021). The attack persisted undetected as users implicitly trusted the tool, with no way to verify its runtime state. Other attacks have also taken this form of abuse (Thomas et al., 2016; Kotzias et al., 2015; Kwon et al., 2017).
Trust Assumptions
We assume legitimate signers do not sign untrusted code, verifiers execute verification procedure correctly, the trusted execution environment setup remains secure, and not all Scope providers can be compromised at the same time. Rationales and standard assumptions are discussed in Appendix A.
4. System Goals and Overview
We now describe our deployment and security goals and provides an overview of DiVerify’s signing and verification flow.
4.1. System Goals
In order to address the attackers described in our Threat Model, DiVerify extends authentication beyond traditional credentials and hardens the security posture of signing clients; and limits reliance on any single identity provider by requiring multiple, independent sources of verification. We formalize these notions into these security goals:
@@ 4.1. System Goals
    Resilience Against Identity Theft: Signing events are allowed only after verification from multiple independent sources, which limits trust assumptions on individual identity providers. This way, if an attacker compromises a Signer’s credentials, the attacker is unable to complete a signing request without a collection of these.
@@
    Resilience Against IdP compromise: In addition to diversity of credentials, a Signer can request multiple identity providers to prove a diverse identity claim. This ensures that a compromised provider cannot impersonate its user unless it has also compromised a threshold number of other independent providers. Unlike S1, which protects against attacker control of a signer’s own credentials, S2 protects against an identity provider issuing fraudulent assertions even when the signer’s credentials are uncompromised.
@@
In addition to the security goals, DiVerify is designed with deployability in mind. We describe in section 10.2 how its design enables integration into current code-signing systems with minimal changes.
4.2. Components
@@
manage and issue claims about a signer’s identity, attributes, or environment context. Each operates within a distinct trust domain, that is, an independently managed authority responsible for issuing and validating its own set of claims about signers. They follow an ad-hoc protocol to verify properties such as control of an email address, possession of a cryptographic key, or ownership of a device and then issues a signed statement attesting to that property. DiVerify does not propose new protocols for scope providers, but builds on existing standards whenever possible. Scope providers must be able to authenticate themselves to the tool and prevent scope replay to ensure that claims are both trustworthy and fresh. When native protocols the required properties, DiVerify “wraps” them using a meta-protocol ( section 5.3).
@@
is a signer-facing component responsible for initiating signing requests. This can be instantiated as a command-line tool or library integrated into developer environments. The client operates within the signer’s local context and has access to the artifact to sign. It serves as the point of integration between the signer and the rest of the DiVerify architecture. Because the client runs in a potentially untrusted setting, we split the functionality of a traditional signing client into two parts: a front-end Client and a privileged Daemon.
@@
is an end-user side component responsible for validating signed artifacts and assessing their trustworthiness. It operates independently of the signing process and has no privileged relationship with the signer.
4.3. System Overview
DiVerify introduces three mechanisms to strengthen identity-based software signing: (1)zero-trust signing architecture, provides verifiers with a DiVerify proof that binds diverse identity and client-trust conditions into a verifiable object (2) a DiVerify policy that gives verifiers explicit and evolvable control over the trust conditions under which signatures are accepted; Figure 2 illustrates how each mechanism integrates into the signing workflow.
These mechanisms are built on a disaggregated trust model that replaces traditional reliance on a single identity provider with a set of independent scope providers. Each scope provider asserts a specific property about the signer such as the identity, the device used, or the presence of another factor. When combined with attestation of the signing client’s integrity, these claims require an attacker to obtain strictly more and qualitatively different capabilities in order to cause a verifier to accept a malicious artifact, compared to signing based on a single identity assertion.
As shown in Figure 2, the signing process begins as in conventional identity-based software signing systems when requests that their client produce a signature over their software artifact. However, instead of obtaining a single identity token from an identity provider, the client quires multiple scope providers to assemble the set of claims required by the verifier’s trust expectations. e.g., a signer may obtain an OIDC token for identity, a fingerprint from the device being used, and a second factor assertion from a hardware token. In parallel, the client collects integrity evidence about its own execution environment.
The client aggregates these claims into a DiVerify proof. The proof captures a diverse security posture of the signing event and depending on the integration path (discussed in section 7), this proof is be distributed alongside the signature or embedded within existing certificate produced by identity-based software signing tools.
On the verifier side, trust decisions are guided by a DiVerify policy, which specifies the scopes, attestation properties, and the logical conditions that must hold for a signature to be trustworthy. e.g., a policy for Alice may require that software be signed by alice@software.sh, from a known device and using a registered hardware second factor. When Bob retrieves a signed artifact, he validates the signature and checks that its its accompanying DiVerify proof satisfies the policy. The policy therefore replaces the implicit trust delegation traditionally provided by identity providers with an explicit specification of the trusted signing conditions.
Figure 2. DiVerify overview. Scope providers assert signer claims that, together with client integrity attestation, form DiVerify proof. Verifiers check this proof against a policy to decide whether to accept the signed artifact.
As an example, suppose Alice maintains a software package and wants verifiers to trust updates only if she signed them as alice@software.sh, from her MacBook Pro, and using her registered Yubikey. She expresses this intent in a DiVerify policy. During signing, the client collects the required claims from the relevant scope providers and incorporates them into the DiVerify proof. When Bob later verifies the package, he checks that the signature is valid and that the accompanying DiVerify proof satisfies Alice’s policy. An attacker who merely obtains Alice’s password or steals her identity token cannot satisfy these additional contextual requirements, the proof fails the policy check and the signature is rejected. In this way, DiVerify prevents attackers from signing on Alice’s behalf even if some authentication factors are compromised.
We defer the discussion of policy distribution, evolution and revocation to section 6.2, which describes how verifiers obtain and ensure they use the right policy.
5. Generating and Verifying Proofs
DiVerify executes a two-phase protocol to sign and verify software. First it asserts and aggregates scope into proof describing security posture when signing was performed. Second, it validates the scopes and the artifact against some expected state.
5.1. Generating DiVerify Proofs
Proof generation proceeds as a single signing session in which the daemon (i) collects required scope claims, (ii) validates them, and (iii) binds them to the signing key and attestation.
DiVerify client advertises the set of scope providers they support. When the DiVerify daemon starts, it establishes communication channels with the client and with each supported scope provider.
The signing protocol, shown in Algorithm 1, begins when the signer initiates a signing request via the client, specifying the artifact identifier (e.g., a filename or container image tag). The client then forwards this request to the DiVerify daemon.
The daemon determines which scopes are required based on the supported providers. These scopes are independent, and may be obtained in parallel. The daemon generates a nonce to bind all subsequent scope responses to the current signing session. For each required scope provider, the daemon executes a scope request meta-protocol.
@@ 5.1. Generating DiVerify Proofs
1:In: Artifact AA
2:Requires: Scope Provider $P$,
@@
4:procedure SignRequest(AA)
5:  $scopes\leftarrow\{\}$
6:  for $p\in providers$ do
7:   $scopes[p]\leftarrow\text{AuthenticateSigner}(p)$
8:  end for
9:  if not ValidScope(scopesscopes) then
10:   return Err
11:  end if
12:  $sk,pk=\text{GenerateKeyPair}()$
13:  $S\leftarrow\text{Sign}_{sk}(A)$
14:  $custom\_data\leftarrow\text{Sign}_{sk}(hash(scopes))$
15:  $quote\leftarrow\text{GetQuote}(custom\_data)$
16:  $dvp\leftarrow\text{MakeDiVerifyProof}(scopes,quote,pk)$
17:  return $S$, dvpdvp
18:end procedure
For each scope request, the daemon sends a request of the form $request(n,claim)$ to the provider. The provider returns a tuple $\langle\mathit{sign(proof)},\ \mathit{proof}\rangle$, where $\mathit{proof}$ is a protocol-specific claim (e.g., an oidc token) that embeds the nonce, and the accompanying signature authenticates the response.
After collecting responses from all required providers, the daemon verifies each response by checking the provider signature and that the embedded nonce matches the current signing session.
If all scopes validate successfully, the daemon generates an ephemeral signing key pair and signs the artifact using the private key. The daemon then signs a hash of the validated scopes with the signing key and embeds the result into the custom data report of the quote produced by the TEE.
Finally, the daemon constructs a DiVerify proof consisting of the validated scopes, the signing public key, and the remote attestation. The daemon returns the artifact signature and the DiVerify proof to the client, which uploads them to the package registry. Embedding a signature over the validated scopes into the attestation quote binds the signing key and scope set to the TEE’s execution context, preventing replay of the attestation
Algorithm 2 Artifact Verification Flow
1:Input: Artifact A, Policy, ArtifactSignature $S$, DiVerify proof dvpdvp
2:Output: Boolean verification result
3:procedure VerifyArtifact(A, S)
4:  quotequote $\leftarrow$ $dvp.\text{quote}$
5:  custom_datacustom\_data $\leftarrow$ $quote.\text{custom\_data}$
6:  pkpk $\leftarrow$ $dvp.\text{pubkey}$
7:  scope_digestscope\_digest $\leftarrow$ $hash(dvp.scopes)$
8:  if not ValidQuote(quotequote) then
9:   return false
10:  end if
11:  if not VerifySignature(scope_digestscope\_digest,custom_datacustom\_data,pkpk) then
12:   return false
@@
14:  if not EvaluatePolicy($dvp.scopes$, policypolicy) then
15:   return false
16:  end if
17:  if not VerifySignature(AA,$S$,pkpk) then
18:   return false
19:  end if
20:  return true
21:end procedure
5.2. Verifying DiVerify Proofs
To verify a signature, a verifier needs the signed artifact, the artifact signature, the DiVerify proof, and a verification policy specifying the expected signer claims and acceptance rules, defined in section 6.
@@ 5.2. Verifying DiVerify Proofs
The verification protocol follows Algorithm 2. The verifier first retrieves the policy. It then extracts the attestation quote, the embedded custom data, public key, and scope claims from the DiVerify proof. The verifier first verifies the integrity of the attestation quote and then checks that the scope digest matches the custom data using the extracted public key. Next, it evaluates the extracted scopes against the policy. Finally, it verifies the artifact signature using the public key. If any check fails, the verifier rejects the artifact.
5.3. Integrating Scope Providers